The daily stock price for International Business Machines (IBM) historically has followed an approximately normal distribution (when adjusting for inflation) with a mean of $131.048 and standard deviation of $2.7756 Approximately 9.44% of days IBM had a stock price greater than what dollar amount?
Answer #1

Mean, µ = 131.048

Standard deviation, σ = 2.7756

P(x > a) = 0.0944

= 1 - P(x < a) = 0.0944

= P(x < a) = 0.9056

Z score at p = 0.9056 using excel = NORM.S.INV(0.9056) = 1.3141

Value of X = µ + z*σ = 131.048 + (1.3141)*2.7756 = 134.6955 = 134.7